Consider a bank where total loans are equal to $800 million and required reserves are $40
million, and where the reserve requirement is 10%.

a) Fill the bank's balance sheet, assuming that it does not hold excess reserves
b) Write the formula for the money multiplier and compute its value. How does it depend on the
reserve ratio? Negatively or positively?
c) Suppose that the Fed injects $80 billion of new reserves into the banking system. Given the
money multiplier computed in b., how large would be the total change in money supply in the
whole banking system?
1
Expert's answer
2016-05-20T09:00:58-0400
Loans are equal to $800 million, required reserves are $40 million, rr = 10%.
a) If the bank does not hold excess reserves, then its assets are $800 million, liabilities are 40/0.1 = $400 million, capital is $400 million.
b) Money multiplier can be calculated using formula mm = 1/rr = 1/0.1 = 10. Money multiplier has inverse (negative) relationship with the reserve ratio.
c) If the Fed injects $80 billion of new reserves into the banking system, then the total change in money supply in the whole banking system will be 80/0.1 = $800 billion.